OpenStudy (anonymous):

sqrt6x-1=sqrt5 can somebody solve this??

OpenStudy (anonymous):

\[\sqrt{6x-1}=\sqrt{5}\]

OpenStudy (anonymous):

\[\sqrt{6x-1}=\sqrt{5}\]

OpenStudy (anonymous):

\[\sqrt{6x-1}^{2} = \sqrt{5}^{2}\] 6x-1 =5 6x=6 x=1

hero (hero):

first step, square both sides